A common mineral in Earth’s crust has the chemical composition 34.55% Mg, 19.96% Si, and 45.49% O by mass. What is its empirical formula?

please show your steps
Answer #1

step1:

no.of moles of Mg = 34.45/24.3 = 1.417

no.of moles of Si = 19.96/28 = 0.713

no.of moles of O = 45.49/16 = 2.843

step2

divide each no. of moles with lowest among these three

Mg = 1.417/0.713 = 1.98 = 2

Si = 0.713/0.713 = 1 =1

O =2.843/0.713   = 3.98 = 4

step3

To obtain emperical formul aassign these values at the suffix of of Mg , Si , O

so Mg2SiO4